Based on market research, a film production company in Ectenia obtains the following

information about the demand and production costs of its new DVD:

π·π‘’π‘šπ‘Žπ‘›π‘‘: 𝑃 = 1000 – 10𝑄

π‘€π‘Žπ‘Ÿπ‘”π‘–π‘›π‘Žπ‘™ 𝑅𝑒𝑣𝑒𝑛𝑒𝑒: 𝑀𝑅 = 1000𝑄 – 10𝑄2

π‘‡π‘œπ‘‘π‘Žπ‘™ πΆπ‘œπ‘ π‘‘: 𝑇𝐢 = 1000 βˆ’ 20𝑄

π‘€π‘Žπ‘Ÿπ‘”π‘–π‘›π‘Žπ‘™ πΆπ‘œπ‘ π‘‘: 𝑀𝐢 = 100 + 10

Where, Q indicates the number of copies sold and P is the price in Ectenian dollars.


1
Expert's answer
2022-03-07T11:14:50-0500

The maximization rule for

were for a competitive firm is to choose the volume of production at which price and marginal revenue equals marginal cost: P = M R = MC.

1000-20Q=100+10Q

30Q=900

Q=30

P=1000-300=700


"Profit=TR-TC=1000Q-10Q^2-(5Q^2+100Q)=1000\times30-10(30)^2-5(30)^2-100\times30=13500"